OpenStudy (anonymous):

Hannah earned $1536 in 8 years on an investment at a 4% simple interest rate. How much was Hannah's investment? Hannah earned $1536 in 8 years on an investment at a 4% simple interest rate. How much was Hannah's investment? @Mathematics

OpenStudy (kropot72):

The amount of interest earned in 1 year = 1536/8 dollars. 1536/8 is 4% of the investment. Therefore 1% of the investment is\[\frac{1536}{8} \div 4=\frac{1536}{8\times 4}\] The investment must be the previous amount multiplied by 100:\[\frac{1536}{8\times 4}\times 100=you\ can\ calculate\]